OpenStudy (anonymous):

Assume that y varies inversely with x. If y = -3 when x = 2.4, what is y when x = 2?

OpenStudy (ddcamp):

If y varies inversely with x, then x × y = c, where c is a constant (It will always be the same) From the first part, we can find that: c = -3×2.4 = -7.2 So just solve for y: y×2=-7.2